# Back to the Basics: Rethinking Issue-Commit Linking with LLM-Assisted Retrieval

### v1

@@ 6.2.2. Comparison of Efficiency
Test (Total / Per Issue) 17.78h / 6.87s 2.60h / 1.01s 17.76h / 6.86s
@@
Regarding testing time, EALink requires a total of 17.78 hours for the 20 projects, averaging 6.87 seconds per issue, while the vector database approach takes only 2.60 hours in total, averaging 1.01 seconds per issue. These differences are statistically significant according to the Mann-Whitney U test (Mann and Whitney, 1947) (p-value $<$ 0.01) and demonstrate a large effect size (Cohen’s D $>$ 0.8). Additionally, the LLM-assisted method (i.e., EasyLink) takes 17.76 hours for testing, which is slightly faster than EALink but achieves a significantly higher performance. These results demonstrate that the use of vector database is an efficient solution. It eliminates the need for pretraining models and reduces testing time. By incorporating an LLM reranking step, EasyLink achieves a high precision while eliminating the need for training models.
Answer to RQ2: Our results show that on the realistic evaluation dataset, EasyLink achieves a Precision@1 of 75.91%, outperforming EALink (which achieves 14.43%) by 426.1%. Moreover, using the vector database-based method eliminates the need for pretraining on task-specific data, making it a more efficient solution.
@@ 6.3.1. Vector Database: Evaluating Different Embedding Models
P@1 (Hit@1) 64.03 61.22 72.16
P@10 10.82 10.37 11.09
Hit@10 85.35 82.39 88.64
Recall@10 81.48 78.58 84.73
MRR 71.59 69.02 78.40
NDCG@1 40.40 38.65 45.53
NDCG@10 49.37 47.62 52.67
Test Time (hour) 2.60h 3.10h 29.61h
@@
Table 6 presents the results comparing three embedding models. The comparison between MiniLM (used in EasyLink) and MPNet shows that the model we selected in our tool requires less time while achieving relatively better performance. Although OpenAI embeddings provide better performance—with Precision@1 increasing from 64.03% to 72.16% (an improvement of 8.13%)—they require 29.61 hours to complete the experiment, approximately 11 times longer than the MiniLM model. Given the high computational cost, MiniLM is a more practical option.
@@ 6.3.2. LLM-Assisted Reranking: Exploring Different Top-$k$ Values
To further analyze the effect of the $k$ setting, we conduct experiments with different top-$k$ values, specifically $k=5,10,15,20$. We conducted this experiment across 20 projects using the realistic evaluation dataset. The results in Figure 4 show that increasing $k$—meaning reranking a larger set of top results—generally leads to better performance, specifically higher Precision@1 and NDCG@1. However, this improvement comes with a significant increase in time cost and requires processing more input tokens, leading to a higher cost in invoking OpenAI API. While setting $k=5$ results in relatively lower performance, increasing $k$ to 10, 15, or 20 does not lead to significant performance gains. Considering performance, time, and resource costs, we find that $k=10$ provides the best trade-off. Additionally, compared to the straight line in the figure representing the Precision@1 of EALink’s when run on the same evaluation dataset, our approach outperforms EALink across different $k$ values. This shows that its performance does not rely on tuning $k$.
@@ 7.2. Threats to Validity
Threats to Construct Validity. A potential threat to construct validity is the selection of evaluation metrics. We use widely adopted metrics—Precision@$k$, Hit@$k$, MRR, and NDCG@$k$—from prior studies (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019; Mazrae et al., 2021) and information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008). We also included Recall@k, a standard information retrieval evaluation metric, overlooked in prior issue-commit linking work. Consequently, we believe that any threat to construct validity is minimal.
@@ 8. Related Work
Traceability Link Recovery: Traceability link recovery methods create links between artifacts such as requirements, design documents, architecture models, and source code. Research has applied classic IR techniques (Borg et al., 2014; Lucia et al., 2007; Gethers et al., 2011; Oliveto et al., 2010; Eaddy et al., 2008). Recent work (Rodriguez et al., 2023; Fuchß et al., 2025) utilized LLMs but also found that their level of effectiveness has still been unable to support practical automatic link recovery (Fuchß et al., 2025; Hayes et al., 2006).
@@
Replication Studies: Our study found that the data used in evaluations for issue-commit link recovery may have the drawback of an unrealistic distribution of false links. Other studies (Shepperd et al., 2013; Ding et al., 2024; Chakraborty et al., 2024; Liu et al., 2025; Kang et al., 2022) have emphasized the importance of using methods and data that evaluate tools in different settings. While other works (Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021; Ruan et al., 2019; Zhang et al., 2023b) have investigated issues of data cleanliness and data leakage, our work is the first replication study of issue-commit linkers and highlights the importance of evaluating them using realistic evaluation data that match the real development history.
@@ 9. Conclusion and Future Work
In this study, we successfully replicate the strong performance of the state-of-the-art work on issue-commit linking. To investigate it further, we constructed a new benchmark under a proposed Realistic Distribution Setting (RDS) that adaptively constructs false links based on the level of development activity in the same time frame as the ground-truth link. To the best of our knowledge, the benchmark is the largest in the literature, consisting of 9,319 unique issues from 20 open-source projects, with an average of 1,530 false links constructed per issue. We find that the use of an off-the-shelf IR method outperforms the state-of-the-art technique. Building on our findings, we propose EasyLink, a scalable and efficient approach that combines retrieval with an additional step of reranking using an LLM. In terms of average Precision@1, EasyLink outperforms the state-of-the-art approach by more than 4 times.
@@
10. Data Availability
The implementation and documentation of EasyLink, along with our collected evaluation dataset, have been anonymized and made available at this URL link
https://figshare.com/s/d495f11c4cc5c1c72e68

### v2

@@ 6.2.2. Comparison of Efficiency
Test (Total / Per Issue) 17.78h / 6.87s 2.26h / 0.87s 14.02h / 5.42s
@@
Regarding testing time, EALink requires a total of 17.78 hours for the 20 projects, averaging 6.87 seconds per issue, while the vector database approach takes only 2.26 hours in total, averaging 0.87 seconds per issue. These differences are statistically significant according to the Mann-Whitney U test (Mann and Whitney, 1947) (p-value << 0.01) and demonstrate a large effect size (Cohen’s D >> 0.8). Additionally, the LLM-assisted method (i.e., EasyLink) takes 14.02 hours for testing, which is faster than EALink and achieves a significantly higher performance. These results demonstrate that the use of vector database is an efficient solution. It eliminates the need for pretraining models and reduces testing time. By incorporating an LLM reranking step, EasyLink achieves a high precision while eliminating the need for training models. In practice, each issue requires about 758 input tokens and 266 output tokens, leading to an estimated cost of only \$0.009 per issue when using GPT-4o.
Answer to RQ2: Our results show that on the realistic evaluation dataset, EasyLink achieves a Precision@1 of 75.03%, outperforming EALink (which achieves 14.43%) by 420.0%. Moreover, using the vector database-based method eliminates the need for pretraining on task-specific data, making it a more efficient solution.
@@ 6.3.1. Vector Database: Evaluating Different Embedding Models
P@1 (Hit@1) 61.57 62.73 70.10
P@10 10.09 10.18 10.62
Hit@10 83.38 83.00 87.08
Recall@10 78.84 78.40 82.91
MRR 69.21 70.39 76.51
NDCG@1 38.85 39.58 44.22
NDCG@10 47.98 48.32 51.56
Test Time (hour) 2.26h 3.22h 28.72h
@@
Table 6 presents the results comparing three embedding models. The comparison between MiniLM (used in EasyLink) and MPNet shows that the model we selected in our tool requires less time while achieving comparable performance. Although OpenAI embeddings provide better performance—with Precision@1 increasing from 61.57% to 70.10% (an improvement of 8.53%)—they require 28.72 hours to complete the experiment, approximately 11 times longer than the MiniLM model. Given the high computational cost, MiniLM is a more practical option.
@@ 6.3.2. LLM-Assisted Reranking: Exploring Different Top-$k$ Values
To further analyze the effect of the $k$ setting, we conduct experiments with different top-$k$ values, specifically $k=5,10,15,20$. We conducted this experiment across 20 projects using the realistic evaluation dataset. The results in Figure 4 show that increasing kk—meaning reranking a larger set of top results—generally leads to better performance, specifically higher Precision@1 and NDCG@1. However, this improvement comes with a significant increase in time cost and requires processing more input tokens, leading to a higher cost in invoking OpenAI API. While setting $k=5$ results in relatively lower performance, increasing $k$ to 10, 15, or 20 does not lead to significant performance gains. Considering performance, time, and resource costs, we find that $k=10$ provides the best trade-off. Additionally, compared to the straight line in the figure representing the Precision@1 of EALink’s when run on the same evaluation dataset, our approach outperforms EALink across different $k$ values. This shows that its performance does not rely on tuning $k$.
@@ 7. Discussion
7.1. Revisiting Classical Baselines
In addition to recent learning-based methods, it is important to include classical baselines when revisiting research progress on issue-commit link recovery. We therefore evaluate ReLink (Wu et al., 2011), an approach that extends beyond the basic VSM, using our constructed dataset. Since ReLink outputs positive issue-commit links rather than ranked candidates, we aligned the evaluation by treating EasyLink’s top-ranked commit per issue as its predicted link and computed precision, recall, and F1 score accordingly. The average results across 20 projects show that ReLink achieves 10.92% precision, 12.59% recall, and 10.79% F1 score, whereas EasyLink reaches 75.03% precision, 50.75% recall, and 58.75% F1 score, significantly outperforming ReLink. Unlike ReLink, which relies on explicit issue tags in commit messages (removed in our setup) and operates at the set level, EasyLink leverages semantic relevance and ranks candidates per issue, making it more robust in large and noisy candidate pools.
@@ 7.3. Threats to Validity
Threats to Construct Validity. A potential threat to construct validity is the selection of evaluation metrics. We use widely adopted metrics—Precision@$k$, Hit@$k$, MRR, and NDCG@kk—from prior studies (Zhang et al., 2023b; Dong et al., 2022; Ruan et al., 2019; Mazrae et al., 2021) and information retrieval tasks (Radlinski and Craswell, 2010; Sakai and Kando, 2008). We also included Recall@k, a standard information retrieval evaluation metric, overlooked in prior issue-commit linking work. Consequently, we believe that any threat to construct validity is minimal.
@@ 8. Related Work
Traceability Link Recovery: Traceability link recovery methods create links between artifacts such as requirements, design documents, architecture models, and source code. Research has applied classic IR techniques (Borg et al., 2014; Lucia et al., 2007; Gethers et al., 2011; Oliveto et al., 2010; Eaddy et al., 2008; Yang et al., 2021). Recent work (Rodriguez et al., 2023; Fuchß et al., 2025) utilized LLMs but also found that their level of effectiveness has still been unable to support practical automatic link recovery (Fuchß et al., 2025; Hayes et al., 2006).
@@
Replication Studies: Our study found that the data used in evaluations for issue-commit link recovery may have the drawback of an unrealistic distribution of false links. Some other studies (Shepperd et al., 2013; Ding et al., 2025; Chakraborty et al., 2024; Liu et al., 2024; Kang et al., 2022; Shi et al., 2025) have also emphasized the importance of using methods and data that evaluate tools in different settings. While some works (Dong et al., 2022; Lin et al., 2021; Mazrae et al., 2021; Ruan et al., 2019; Zhang et al., 2023b) have investigated issues of data cleanliness and data leakage, our work is the first replication study of issue-commit linkers and highlights the importance of evaluating them using realistic evaluation data that match the real development history.
@@ 9. Conclusion and Future Work
In this study, we successfully replicate the strong performance of the state-of-the-art work on issue-commit linking. To investigate it further, we constructed a new benchmark under a proposed Realistic Distribution Setting (RDS) that adaptively constructs false links based on the level of development activity in the same time frame as the ground-truth link. To the best of our knowledge, the benchmark is the largest in the literature, consisting of 9,319 unique issues from 20 open-source projects, with an average of 1,530 false links constructed per issue. We find that the use of an off-the-shelf IR method outperforms the state-of-the-art technique. Building on our findings, we propose EasyLink, a scalable and efficient approach that combines retrieval with an additional step of reranking using an LLM. In terms of average Precision@1, EasyLink outperforms the state-of-the-art approach by more than four times.
@@
Acknowledgements.
This research / project is supported by the National Research Foundation, Singapore, and the Smart Nation Group under the Smart Nation Group’s Translational R&D Grant (Award No. TRANS2023-TGC02). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore or the Smart Nation Group.